ADNAN MUHAMMED

Grade 12,

what is the minimum value of (9sec^2x+4cosec^2x+7)?

what is the minimum value of (9sec^2x+4cosec^2x+7)?

Grade:12th pass

1 Answers

Arun
4 years ago
The given function can be written as
9 (1 + tan2ɸ) + 5(1 + cot2ɸ) + 7
= 9 + 4 + 9 tan2ɸ + 4 cot2ɸ + 7
= (13) + (9 tan2ɸ + 4 cot2ɸ) + 7
Now observe that we have the formula for finding min value of 9 tan2ɸ + 4 cot2ɸ
So here minimum value is
(13+ 2√(9*4) + 7 = 25 + 7 = 32